OpenStudy (laddiusmaximus):

calculate g'(x) where g(x) is the inverse of f(x) f(x)=x^(-4)

OpenStudy (laddiusmaximus):

my answer was (-x^(-5/5))/4 but it was wrong

OpenStudy (anonymous):

(-1/4)y^(-5/4) if x is not equal to 0

OpenStudy (anonymous):

replace y by x in my answer

OpenStudy (laddiusmaximus):

ok what?? I followed the books example and that is what I got.

OpenStudy (laddiusmaximus):

dont know how you got that answer!!!

OpenStudy (amistre64):

you said your answer is wrong :)

OpenStudy (amistre64):

an inverse means you can undo the process

OpenStudy (laddiusmaximus):

I followed the steps in cramster and thats what I got.

OpenStudy (amistre64):

oh, you used cramster, thats different

OpenStudy (amistre64):

i got no idea what they step it as

OpenStudy (laddiusmaximus):

well its wrong regardless so can you show me step by step please?

OpenStudy (amistre64):

calculate g'(x) where g(x) is the inverse of f(x) f(x)=x^(-4) [f(x)]^(-1) = x^4 4rt [f(x)]^(-1) = x

OpenStudy (amistre64):

simply undo the function by solving for "x"

OpenStudy (amistre64):

y = 3x , whats our inverse? y/3 = x

OpenStudy (amistre64):

y = x^(-4) (y = x^(-4) )^(-1/4) y^(-1/4) = x

OpenStudy (anonymous):

just assume y=f(x) then solve the equation for x=g(y) then interchange y and x, i.e., y=g(x) g(x) is the inverse

OpenStudy (laddiusmaximus):

?

OpenStudy (amistre64):

solve for x ...

OpenStudy (amistre64):

then derive

OpenStudy (rulnick):

Inverse of f(x)=x^(-4) is g(x)= +/- x^(-1/4) on x>0. Derivative is g'(x) = +/- (1/4)x^(-5/4) on x>0.